# astroCAMP: A Community Benchmark and Co-Design Framework for Sustainable SKA-Scale Radio Imaging

### v1

@@ 3.2. Co-design Metrics
Table 2. astroCAMP co-design metrics. Core metrics (C) enter the performance–energy–quality optimisation; diagnostic metrics (D) support interpretation and reproducibility. Type symbols: Direct = $\bullet$, Derived = $\circ$, Proxy/Model-based = $\triangle$.
@@
Layer ID Metric / Formula Unit Role Type Instrumentation Interpretation
@@
System (Pipeline) A1 Time-to-solution $T_{c}$ s C $\bullet$ POSIX time, workflow logs, scheduler timestamps End-to-end wall-clock runtime of the workload on the full (possibly heterogeneous) system, where $T_{c}$ is the elapsed execution time.
@@
A2 Energy-to-solution $E_{c}=\int_{0}^{T_{c}}P(t)\,dt$ J C $\bullet$ Rack/node PDUs, CPU RAPL, GPU NVML, PMT traces Total electrical energy consumed by the system, integrating instantaneous power $P(t)$ over runtime $T_{c}$.
@@
A3 Throughput $\Theta=N/T_{c}$ vis/s C $\circ$ MS/FITS logs, visibility/image counters System-level processing rate, where $N$ is the number of visibilities or images processed during runtime $T_{c}$.
@@
A4 Energy efficiency $\eta_{E}=N/E_{c}$ vis/J C $\circ$ Derived from $N$ and $E_{c}$ System-level science throughput per joule, using processed data volume $N$ and energy-to-solution $E_{c}$.
@@
Hardware Platform A5 Utilisation $U=t_{\mathrm{active}}/t_{\mathrm{total}}$ C $\bullet$ nvidia-smi, ROCm tools, perf, Prometheus exporters Fraction of time a device (CPU socket, GPU, FPGA, etc.) is actively running kernels, where $t_{\mathrm{active}}$ is active compute time and $t_{\mathrm{total}}$ is total wall time.
@@
A6 Memory bandwidth $B_{\mathrm{mem}}=\mathrm{Bytes}/T_{c}$ GB/s D $\bullet$ Hardware counters, Intel VTune, NVIDIA Nsight Sustained data-movement rate per platform, with $B_{\mathrm{mem}}$ indicating whether kernels are memory-bound.
@@
A7 Peak memory usage $M_{\mathrm{peak}}$ GB C $\bullet$ /proc/meminfo, nvidia-smi, cgroups, container telemetry Maximum resident memory footprint observed on a platform during execution, constraining batch size and problem scaling.
@@
Algorithmic Quality B1 Dirty-image RMS $\sigma_{\mathrm{dirty}}=\sqrt{\frac{1}{N}\sum(I_{i}-\bar{I})^{2}}$ Jy/beam C $\bullet$ CASA/WSClean imstat, PyBDSF Noise and artefact level in the dirty image, using pixel intensities $I_{i}$, mean value $\bar{I}$, and resulting RMS $\sigma_{\mathrm{dirty}}$.
@@
B2 PSNR / SSIM PSNR $10\log_{10}(I_{\max}^{2}/\mathrm{MSE})$ dB / – C $\bullet$ scikit-image, OpenCV Fidelity of reconstruction $\hat{I}$ vs. reference $I_{\mathrm{ref}}$, using maximum pixel $I_{\max}$ and mean-squared error; SSIM measures structural similarity between $\hat{I}$ and $I_{\mathrm{ref}}$.
@@
B3 Dynamic range $DR=I_{\max}/\sigma_{\mathrm{res}}$ C $\bullet$ Residual-image statistics Ratio of peak intensity $I_{\max}$ to residual RMS $\sigma_{\mathrm{res}}$, determining detectability of faint emission.
@@
B4 Astrometric error $\epsilon_{\mathrm{astro}}=\frac{1}{N}\sum\|\mathbf{x}_{i}-\mathbf{x}_{i}^{\mathrm{ref}}\|$ arcsec or px C $\bullet$ PyBDSF catalogues Average positional discrepancy between reconstructed and reference sources, where $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{\mathrm{ref}}$ are measured and reference positions.
@@
Sustain. C1 Carbon-to-solution $C_{c}=E_{c}\,\kappa(t,r)$ gCO2e C $\triangle$ Grid carbon-intensity APIs (e.g. electricityMap) + $E_{c}$ Total carbon footprint of a run, using measured energy $E_{c}$ and grid carbon intensity $\kappa(t,r)$ at time $t$ and region $r$.
@@
C2 Carbon efficiency $\eta_{C}=N/C_{c}$ vis/gCO2e D $\circ$ Derived from $N$ and $C_{c}$ Visibilities processed per gram CO2 emitted, normalising science throughput $N$ by carbon cost $C_{c}$.
@@
Economic E1 Total cost ownership $C_{\mathrm{TTO}}=C_{\mathrm{capex}}+C_{\mathrm{opex}}$ D $\bullet$ Procurement & operations records Lifetime system cost, combining capital cost $C_{\mathrm{capex}}$ and operational cost $C_{\mathrm{opex}}$.
@@
E2 Cost per job $C_{E}=E_{c}\,p_{E}$ C $\circ$ Electricity tariff + $E_{c}$ Monetary cost per workload, using energy-to-solution $E_{c}$ and electricity price $p_{E}$.
@@
E3 Cost efficiency $\Theta/C_{\mathrm{TTO}}$ vis/\$ D $\circ$ Throughput + TTO data Operations delivered per euro invested, using throughput $\Theta$ and total cost $C_{\mathrm{TTO}}$.
@@
astroCAMP includes a compact, cross-layer metric suite that links hardware behaviour, energy use, economic cost, and scientific fidelity. The framework measures hardware architecture (platform) and system performance using time-to-solution $T_{c}$, energy-to-solution $E_{c}$, average power, sustained memory bandwidth, utilisation, and domain-specific throughput (visibilities/s or images/s). Algorithmic quality is quantified via dirty-image RMS, PSNR/SSIM, dynamic range, and astrometric and photometric errors from catalogue comparisons, capturing the scientific impact of approximations such as reduced precision, coarse $w$-stacking, or NUFFT kernel truncation. Sustainability metrics pair energy use with regional carbon intensity $\kappa(t,r)$ to compute Carbon-to-solution $C_{c}$, enabling joint optimisation of energy, carbon, and image fidelity. Economic metrics—including cost per job $C_{E}=E_{c}\,p_{E}$ and amortised total cost of ownership $C_{\mathrm{TTO}}$—connect computational choices to financial constraints relevant for large-scale SKA deployments.
Together, the full metric suite supports transparent, multi-objective co-design across CPU, GPU, FPGA, and ASIC platforms. Table 2 summarises all metrics, including their formulas, units, and instrumentation sources.
@@ 4. Extensible Multi-Objective Co-Design Formulation
Each $x\in\mathcal{P}$ represents a distinct, scientifically valid, and operationally feasible trade-off (e.g., “fast but power-hungry”, “high precision but slower”, “energy-efficient but low throughput”). The construction of $\mathcal{P}$ is done via structured parameter sweeps at the software and/or the hardware architecture level. Evaluation of co-design metrics can be performed using simulation tools (sarkar_lightningsim_2023), model-based design (PREESM2017) or surrogate models. Automated design space exploration (zaourar2023deca) can rely on Bayesian optimization (fu2025high), or evolutionary search to analyze how both new (BIPP, WS-Snapshot) and legacy algorithms (IDG) populate the multi-metric design space.
@@
The datasets were generated using OSKAR (OSKAR) (version OSKAR-2.11.2-dev 2025-07-17 ddb65ed) using SKA-Low 512 stations full configuration. Sixteen datasets were produced varying the number of timesteps (or time samples) over 1, 8, 64, 128, and 256, and the number of channels over 1, 8, 64, 128, and 256. Each timestep is integrated over 10 seconds. Phase center was set to 25.0 and -30.0 degrees in right ascension and declination, respectively. Start frequency was set to 151 MHz with increments of 1 MHz. Sources were selected from the GLEAM catalog (gleam2017) to cover a field of view of 40 degrees. Data volumes range from megabytes to terabytes, enabling tests from workstation to cluster scale. Each dataset includes:
@@ 6. Benchmark Evaluation and Results
This section benchmarks the proposed SKA imaging workloads and datasets using SKA-Low configurations to demonstrate the practical impact of our co-design methodology. We present the experimental setup used to obtain reproducible measurements of a subset of the proposed metrics for system-level performance and energy efficiency, sustainability, and economy. We then evaluate heterogeneous WSClean+IDG deployments and analyse the weak scaling behaviour of CPU-only execution. Assuming we were to run identical workloads at the two SKA SDP locations, we quantify how site-specific carbon intensity and electricity prices shape cost and carbon efficiency. Finally, we use PREESM to explore the co-design space for performance–energy-utilization trade-offs. This demonstrates how the proposed methodology can aid in identifying Pareto-optimal operating regions for SKA-scale imaging pipelines.
@@ 6.1. Experimental Setup
Benchmarks were executed on a Lenovo ThinkSystem SR675 V3 node of the Green500 KUMA cluster (kuma_node_spec), equipped with dual AMD EPYC 9334 CPUs, 371 GB RAM, 6.4 TB NVMe storage, and four NVIDIA H100 GPUs (94 GB). All runs used exclusive node access, with tests using 16 CPU cores, a single H100 GPU and a quarter (92.2 GB) of the full RAM (Slurm configuration limits the memory per CPU core to 5900 MB).
The server cost was obtained by configuring a system with comparable specifications, resulting in $29\,875.39\text{\,}\mathrm{U}\mathrm{S}\mathrm{D}$ (lenovo_sr675_v3). Embodied hardware emissions were estimated by combining the emissions reported by NVIDIA for the H100 GPU ( (nvidia_h100_pcf)) with the emissions of the remaining system components modeled using Boavizta’s Datavizta tool (boavizta_datavizta), yielding a total embodied emission of . The latter assumes a Genoa-class CPU architecture, Samsung memory modules, and a Micron SSD. We assume that the CAPEX is amortized over a six-year lifetime.
The data were processed using WSClean (offringa2014wsclean) using IDG (vandertol2018idg) gridder GPU mode. For each of the 16 datasets we produced images with fixed FoV but varying image sizes of $4096^{2}$, $8192^{2}$, $16\,384^{2}$ and $32\,768^{2}$ pixels, i.e. with pixel scales of respectively 17.578, 8.789, 4.394 and 2.197 arcsec. For each of the $25\times 4=100$ runs, the data was processed twice in the following sequence: warmup, 120 s pause, run, 120 s pause. The warmup run is not monitored but GPU CUDA kernels of IDG are being compiled while the run is monitored and uses the pre-compiled GPU kernels from warmup stage. The 120 s pause before the monitored run allows the hardware to settle down. Power measurements were recorded using node-level PDUs, which provided a 1 s power average sampled every 5 s. The PDU measurements were cross-validated with software monitors (RAPL, NVML), using the PMT library wrapper (corda2022pmt).
@@ 6.2. Heterogeneous WSClean+IDG Deployment
WSClean+IDG executes as a heterogeneous pipeline in which the CPU handles visibility batching, metadata preparation, and kernel scheduling, while the GPU performs all IDG gridding and interpolation. This division achieves high throughput only when the CPU can supply work at a rate sufficient to keep the GPU busy.
Figure 5. Performance metrics of one run of WSClean+IDG.
@@
Fig. 5 shows the measurements for one experiment, including CPU utilization, memory footprint and I/O intensity; CPU, GPU and total system power. Power is plotted as a delta over the power measured during idle period. First, the Fig. shows that, most of the time, CPU utilization is limited to one core (100 %), with short peaks of multi-core execution. This reflects the low scalability of the algorithms on the number of CPU cores. Second, GPU power is on average much lower than the rated TDP of the H100, and the utilization of the GPU’s streaming multiprocessors (SMs) is very low, with just short bursts of activity; this shows that either the CPUs are not able to generate work for the GPU promptly, or the GPU is bottlenecked by the access to system memory.
Fig. 6 reveals a clear energy and performance hierarchy for the largest image size ($32768^{2}$) across all $(n_{\mathrm{times}},n_{\mathrm{chans}})$ configurations: total consumption is overwhelmingly dominated by the GPU and IDG device, which together account for roughly 60–90% of system energy, while CPU, IDG host, and system-level overheads remain small, flat, and secondary contributors. As workloads scale, GPU energy increases sharply (from $\sim$5–10,kJ for the smallest cases to $\sim$165,kJ for t256t256c256c256). In contrast, CPU and host-side IDG energy stay nearly constant, highlighting the accelerator-driven nature of large-scale imaging. Execution time and energy scale together over almost four orders of magnitude, with the longest runtimes corresponding to the highest energy use, and tightly coupled IDG host/device times indicating a well-balanced heterogeneous execution path. Finally, the comparison across configurations shows that channel-heavy workloads achieve similar throughput with lower wall time and total energy than time-heavy cases, making them systematically more energy-efficient.
@@
We analyze with Fig. 7 how energy and throughput vary across the astroCAMP dataset configurations. As image size and visibility dimensions increase, throughput improves initially but rapidly saturates, while wall time and total energy grow super-linearly. This behaviour indicates a transition from a compute-efficient regime to a memory- and bandwidth-bound execution regime, in which additional work no longer translates into proportional gains in Mvis/h. Consequently, naïvely scaling imaging parameters increases operational cost and carbon footprint without delivering commensurate scientific throughput, underscoring the need for efficiency-aware configuration selection for SKA-scale deployments.
6.3. Weak CPU-only Scaling of WSClean+IDG

### v2

@@ 3.2. Co-design Metrics
Table 2. astroCAMP co-design metrics. Core metrics (C) enter the performance–energy–quality-cost optimisation; diagnostic metrics (D) support interpretation and reproducibility. Type symbols: Direct = $\bullet$, Derived = $\circ$, Proxy/Model-based = $\triangle$.
@@
Layer Metric / Formula Unit Role Type Instrumentation Interpretation
@@
System (Pipeline) Time-to-solution $T_{c}$ s C $\bullet$ POSIX time, workflow logs, scheduler timestamps End-to-end wall-clock runtime of the workload on the full (possibly heterogeneous) system, where $T_{c}$ is the elapsed execution time.
@@
Energy-to-solution $E_{c}=\int_{0}^{T_{c}}P(t)\,dt$ J C $\bullet$ Rack/node PDUs, CPU RAPL, GPU NVML, PMT traces Total electrical energy consumed by the system, integrating instantaneous power $P(t)$ over runtime $T_{c}$.
@@
Throughput $\Theta=N/T_{c}$ vis/s C $\circ$ MS/FITS logs, visibility/image counters System-level processing rate, where $N$ is the number of visibilities or images processed during runtime $T_{c}$.
@@
Energy efficiency $\eta_{E}=N/E_{c}$ vis/J C $\circ$ Derived from $N$ and $E_{c}$ System-level science throughput per joule, using processed data volume $N$ and energy-to-solution $E_{c}$.
Data locality $L_{d}=V_{\mathrm{local}}/V_{\mathrm{total}}$ D $\triangle$ Darshan I/O profiler, filesystem telemetry Fraction of total I/O volume served locally, where $V_{\mathrm{local}}$ and $V_{\mathrm{total}}$ are local and total I/O volumes.
@@
Hardware Platform Utilisation $U=t_{\mathrm{active}}/t_{\mathrm{total}}$ D $\bullet$ nvidia-smi, ROCm tools, perf, Prometheus exporters Fraction of time a device (CPU socket, GPU, FPGA, etc.) is actively running kernels, where $t_{\mathrm{active}}$ is active compute time and $t_{\mathrm{total}}$ is total wall time.
@@
Memory bandwidth $B_{\mathrm{mem}}=\mathrm{Bytes}/T_{c}$ GB/s D $\bullet$ Hardware counters, Intel VTune, NVIDIA Nsight Sustained data-movement rate per platform, with $B_{\mathrm{mem}}$ indicating whether kernels are memory-bound.
Memory efficiency $\eta_{\mathrm{mem}}=\mathrm{Bytes}/E_{c}$ GB/J D $\bullet$ Derived from memory-traffic counters and energy-to-solution $E_{c}$ Useful data movement per joule at system level, computed as total DRAM bytes transferred divided by system energy-to-solution $E_{c}$. Complements $B_{\mathrm{mem}}$ by quantifying the energy cost of memory traffic.
@@
Peak memory usage $M_{\mathrm{peak}}$ GB D $\bullet$ /proc/meminfo, nvidia-smi, cgroups, container telemetry Maximum resident memory footprint observed on a platform during execution, constraining batch size and problem scaling.
@@
Algorithmic Quality Dirty-image RMS $\sigma_{\mathrm{dirty}}=\sqrt{\frac{1}{N}\sum(I_{i}-\bar{I})^{2}}$ Jy/beam C $\bullet$ CASA/WSClean imstat, PyBDSF Noise and artefact level in the dirty image, using pixel intensities $I_{i}$, mean value $\bar{I}$, and resulting RMS $\sigma_{\mathrm{dirty}}$.
@@
PSNR / SSIM, PSNR $10\log_{10}(I_{\max}/\mathrm{MSE})$ dB / – C $\bullet$ scikit-image, OpenCV Fidelity of reconstruction $\hat{I}$ vs. reference $I_{\mathrm{ref}}$, using maximum pixel $I_{\max}$ and mean-squared error; SSIM measures structural similarity between $\hat{I}$ and $I_{\mathrm{ref}}$.
@@
Dynamic range $DR=I_{\max}/\sigma_{\mathrm{res}}$ C $\bullet$ Residual-image statistics Ratio of peak intensity $I_{\max}$ to residual RMS $\sigma_{\mathrm{res}}$, determining detectability of faint emission.
@@
Astrometric error $\epsilon_{\mathrm{astro}}=\frac{1}{N}\sum\|\mathbf{x}_{i}-\mathbf{x}_{i}^{\mathrm{ref}}\|$ arcsec or px C $\bullet$ PyBDSF catalogues Average positional discrepancy between reconstructed and reference sources, where $\mathbf{x}_{i}$ and $\mathbf{x}_{i}^{\mathrm{ref}}$ are measured and reference positions.
@@
Sustain. Carbon-to-solution $C_{c}=E_{c}\,\kappa(t,r)$ gCO2e C $\triangle$ Grid carbon-intensity APIs (e.g. electricityMap) + $E_{c}$ Total carbon footprint of a run, using measured energy $E_{c}$ and grid carbon intensity $\kappa(t,r)$ at time $t$ and region $r$.
@@
Carbon efficiency $\eta_{C}=N/C_{c}$ vis/gCO2e D $\circ$ Derived from $N$ and $C_{c}$ Visibilities processed per gram CO2 emitted, normalising science throughput $N$ by carbon cost $C_{c}$.
@@
Economic Total cost ownership $C_{\mathrm{TCO}}=C_{\mathrm{capex}}+C_{\mathrm{opex}}$ D $\bullet$ Procurement & operations records Lifetime system cost, combining capital cost $C_{\mathrm{capex}}$ and operational cost $C_{\mathrm{opex}}$.
@@
Cost per job $C_{E}=E_{c}\,p_{E}$ C $\circ$ Electricity tariff + $E_{c}$ Monetary cost per workload, using energy-to-solution $E_{c}$ and electricity price $p_{E}$.
@@
Cost efficiency $\Theta/C_{\mathrm{TCO}}$ vis/€ D $\circ$ Throughput + TTO data Science operations delivered per euro invested, using throughput $\Theta$ and total cost $C_{\mathrm{TCO}}$.
astroCAMP includes a compact, cross-layer metric suite that links hardware behaviour, energy use, economic cost, scientific throughput, and fidelity. Table 2 summarises all metrics, including their formulas, units, role, and instrumentation, and interpretation. The objective function is to minimize time-, energy-, carbon, and cost-to-solution while maximizing science throughput. Only the core metrics (C) are target for optimization, while diagnostic metrics (D) support the interpretation of results by revealing system-level and platform-level bottlenecks—such as remote data access, uneven work distribution across CPUs or GPUs, and delays caused by slower devices.
@@
The framework captures system-level (pipeline) key performance indicators, including time-to-solution $T_{c}$, energy-to-solution $E_{c}$, and scientific throughput (visibilities/s or images/s), as well as hardware platform (architecture) metrics to inform co-design decisions, such memory bandwidth, memory efficiency, and utilisation. Algorithmic quality is quantified via dirty-image RMS, PSNR/SSIM, dynamic range, and astrometric and photometric errors from catalogue comparisons, capturing the scientific impact of approximations such as reduced precision, coarse $w$-stacking, or NUFFT kernel truncation. Sustainability metrics pair energy use with regional carbon intensity $\kappa(t,r)$ to compute carbon-to-solution $C_{c}$, enabling joint optimisation of energy, carbon, and image fidelity. Economic metrics—including cost per job $C_{E}=E_{c}\,p_{E}$ and amortised total cost of ownership $C_{\mathrm{TCO}}$—connect computational choices to financial constraints relevant for large-scale SKA deployments. Together, the full metric suite supports transparent, multi-objective co-design across CPU, GPU, FPGA, and ASIC platforms.
@@ 4. Extensible Multi-Objective Co-Design Formulation
Each $x\in\mathcal{P}$ represents a distinct, scientifically valid, and operationally feasible trade-off (e.g., “fast but power-hungry”, “high precision but slower”, “energy-efficient but low throughput”). The construction of $\mathcal{P}$ is done via structured parameter sweeps at the software and/or the hardware architecture level. Evaluation of co-design metrics can be performed using simulation tools (Sarkar and Hao, 2023), model-based design (Suriano et al., 2017) or surrogate models. Automated design space exploration (Zaourar et al., 2023) can rely on Bayesian optimization (Fu et al., 2025), or evolutionary search to analyze how both new (BIPP, WS-Snapshot) and legacy algorithms (IDG) populate the multi-metric design space.
@@
The datasets were generated using OSKAR (Mort et al., 2010) (version OSKAR-2.11.2-dev 2025-07-17 ddb65ed) using SKA-Low 512 stations full configuration. Sixteen datasets were produced varying the number of timesteps (or time samples) over 1, 8, 64, 128, and 256, and the number of channels over 1, 8, 64, 128, and 256. Each timestep is integrated over 10 seconds. Phase center was set to 25.0 and -30.0 degrees in right ascension and declination, respectively. Start frequency was set to 151 MHz with increments of 1 MHz. Sources were selected from the GLEAM catalog (Hurley-Walker et al., 2017) to cover a field of view of 40 degrees. Data volumes range from megabytes to terabytes, enabling tests from workstation to cluster scale. Each dataset includes:
@@ 6. Benchmark Evaluation and Results
This section benchmarks SKA-Low workloads and datasets to demonstrate the practical impact of our co-design methodology. We report results from a subset of metrics (performance, energy efficiency, sustainability, and economy), then analyse heterogeneous WSClean+IDG execution, CPU-only strong scaling, location-dependent carbon/cost efficiency for identical workloads, and PREESM-based design-space exploration for performance–energy–utilization trade-offs toward Pareto-optimal SKA-scale operating regions.
@@ 6.1. Experimental Setup
Benchmarks were executed on a Lenovo ThinkSystem SR675 V3 node of the Green500 KUMA cluster (SCITAS, ), with dual AMD EPYC 9334 CPUs, 371 GB RAM, 6.4 TB NVMe, and four NVIDIA H100 GPUs (94 GB). All runs used exclusive node access and allocated 16 CPU cores, one H100 GPU, and one quarter of RAM (92.2 GB; Slurm limit 5900 MB/core).
The server cost was obtained by configuring a system with comparable specifications, resulting in $29\,875.39\text{\,}\mathrm{U}\mathrm{S}\mathrm{D}$ (Lenovo, 2024). Embodied hardware emissions were estimated by combining the emissions reported by NVIDIA for the H100 GPU ( (NVIDIA, 2023)) with the emissions of the remaining system components modeled using Boavizta’s Datavizta tool (Boavizta, 2024), yielding a total embodied emission of . The latter assumes a Genoa-class CPU architecture, Samsung memory modules, and a Micron SSD. We assume that the CAPEX is amortized over a six-year lifetime.
Figure 5. Profiling of WSClean + IDG on GPU: 256 timesteps, 256 channels, full SKA-Low, for a 16k $\times$ 16k pixels image.
The data were processed with WSClean (Offringa et al., 2014) in IDG (Van der Tol et al., 2018a) GPU mode. For each of 16 datasets, we fixed FoV and varied image size ($4096^{2}$, $8192^{2}$, $16\,384^{2}$, $32\,768^{2}$; pixel scales 17.578, 8.789, 4.394, 2.197 arcsec). For each of the $25\times 4=100$ runs, we executed: warmup, 120 s pause, monitored run, 120 s pause. Warmup compiles CUDA kernels; the monitored run reuses compiled kernels. The pre-run pause allows hardware to settle. Power measurements were recorded using node-level PDUs, which provided a 1 s power average sampled every 5 s. The PDU measurements were cross-validated with software monitors (RAPL, NVML), using the PMT library wrapper (Corda et al., 2022a). Fig. 5 exemplifies CPU utilization, memory footprint, I/O intensity, and CPU/GPU/system power (as deltas from idle). CPU activity is mostly confined to one core (100 %) with short multi-core bursts, indicating poor CPU-side scaling. GPU power remains well below H100 TDP and streaming multiprocessors (SM) utilization appears in short bursts, indicating that work supply is intermittent and/or constrained by memory-access effects.
@@ 6.2. Heterogeneous WSClean+IDG Deployment
Figure 8. Memory-efficiency and footprint trends. Lines report memory efficiency for GPU vs CPU execution of IDG.
Figure 9. Largest-image saturation signals grouped by channels $c$. Left: summed explicit IDG sub-kernel times (gridder, sub-FFT, w-tiling) versus end-to-end WSClean wall time for each $(t,c)$ workload. Right: throughput with active-window GPU utilization and memory-footprint statistics (averaged only over samples with active GPU work).
Figure 10. GPU-vs-CPU roofline comparison for WSClean stacking ($t=256$, $c=256$, $16384^{2}$). The figure contrasts arithmetic intensity, achieved throughput, and bandwidth context for CPU-only execution and the GPU-enabled kernels. Roofline markers shape identifies the workload and color the parallelism.
Figure 11. CPU-only roofline context for WSClean stacking ($t=256$, $c=256$, $16384^{2}$) across $p=\{1,16,32,64\}$: arithmetic intensity, achieved throughput, achieved memory bandwidth, and machine ceilings.
WSClean+IDG is a heterogeneous pipeline: the CPU performs batching, metadata preparation, and scheduling, while the GPU executes IDG gridding and interpolation. Throughput is high only when the CPU feeds work fast enough to keep the GPU busy. Fig. 6 shows a clear hierarchy for $32768^{2}$ across $(n_{\mathrm{times}},n_{\mathrm{chans}})$: GPU and IDG-device energy dominate (about 60–90%), while CPU, IDG-host, and system overheads are smaller and comparatively flat. GPU energy rises from about 5–10 kJ to about 165 kJ at t256t256c256c256, whereas CPU/host-side IDG energy remains nearly constant, confirming accelerator-driven scaling. Execution time and energy grow together over nearly four orders of magnitude, and channel-heavy workloads deliver similar throughput with lower wall time and total energy than time-heavy cases.
@@
We analyze with Fig. 7 how energy and throughput vary across the astroCAMP dataset configurations. As visibility dimensions increase, throughput improves initially but rapidly saturates, while wall time and total energy grow super-linearly. This trend reflects increasing data-movement and orchestration overheads for larger working sets, rather than a change in the fundamental computational bottleneck. Consequently, naïvely scaling imaging parameters increases operational cost and carbon footprint without delivering commensurate scientific throughput, underscoring the need for efficiency-aware configuration selection for SKA-scale deployments. Darshan profiling shows that I/O contributes only a small share of total runtime ($\sim 1\%$) for this workload, but is composed of many short operations. For the largest case, we also observe many sub-megabyte reads and a high fraction of misaligned requests, indicating that access granularity and request alignment are relevant optimization targets even when peak filesystem bandwidth is not saturated. Data-movement efficiency metric (Bytes/J) quantifies the energetic cost of memory traffic. Complementing this, Fig. 8 provides an analytical decomposition of transferred workload size and shows that scaling is increasingly input-dominated at larger $(t,c)$, which motivates optimization of ingest, staging, reuse, and data-format/compression choices.
Fig. 8 shows payload-footprint and memory-efficiency trends at $32768^{2}$: increasing channels mainly grows input payload, while the output term remains fixed per image size (0.0625 GiB for $4096^{2}$, 0.25 GiB for $8192^{2}$, 1.0 GiB for $16384^{2}$, and 4.0 GiB for $32768^{2}$). The GB/J curves indicate that larger spectral workloads do not translate into proportional memory-efficiency gains; the device path is generally above host, but both are sublinear versus footprint growth. The co-design implication is to prioritize channel-aware batching, fewer redundant visibility passes, lower orchestration overhead, and scheduling that keeps efficient GPU phases dominant for longer.
The analysis in Fig. 9 shows that increasing $c$ raises explicit IDG kernel work, but end-to-end WSClean wall time remains much larger than summed sub-kernel times, so application-level gains are not proportional. Throughput flattens before active-window SM utilization saturates and while active memory utilization stays modest, indicating a workflow-level bottleneck. This matches Nsight Compute: some kernels are efficient while active, but those states are not sustained across the full workflow timeline, so gains are more likely from batching, overlap, synchronization reduction, and staged-data reuse tha n from kernel-only tuning.
The roofline study in Figs. 10 and 11 adds model-guided context for CPU-only versus GPU-enabled execution. The CPU-only path remains to the right of the DRAM ridge point for this workload while achieving only a modest fraction of peak compute, and the GPU-enabled path shows efficient kernels only in intermittent windows. Together, the end-to-end limits are better explained by orchestration, synchronization, and host-device pipeline effects than by a single DRAM-roof constraint.
6.3. Strong CPU-only Scaling Limitations
[…]